Explain Momentum and its associated principles.


1
Expert's answer
2021-05-06T18:47:20-0400

The momentum of mechanical body is a product of its velocity and mass.

"p=mv"

The unit of momentum in SI is "kg \cdot m/s^2" .

If there are no external forces acting on the system of bodies, then the total momentum is conserved

"p_1 +p_2 + ...+p_N = const"

The second Netwon's law can be written in terms of momentum:

"\displaystyle \frac{dp}{dt} = F"